# When People are Floods: Analyzing Dehumanizing Metaphors in Immigration Discourse with Large Language Models

### v1

@@ Abstract
Metaphor, discussing one concept in terms of another, is abundant in politics and can shape how people understand important issues. We develop a computational approach to measure metaphorical language, focusing on immigration discourse on social media. Grounded in qualitative social science research, we identify seven concepts evoked in immigration discourse (e.g. water or vermin). We propose and evaluate a novel technique that leverages both word-level and document-level signals to measure metaphor with respect to these concepts. We then study the relationship between metaphor, political ideology, and user engagement in 400K US tweets about immigration. While conservatives tend to use dehumanizing metaphors more than liberals, this effect varies widely across concepts. Moreover, creature-related metaphor is associated with more retweets, especially for liberal authors. Our work highlights the potential for computational methods to complement qualitative approaches in understanding subtle and implicit language in political discourse.
@@
Julia Mendelsohn University of Chicago juliame@uchicago.edu                        Ceren Budak University of Michigan cbudak@umich.edu
@@ 1 Introduction
Metaphor, communication of one concept in terms of another, is abundant in political discourse. Metaphor structures how we understand the world (Lakoff and Johnson, 1980), and is deployed consciously and subconsciously to structure our understanding of political issues in terms of accessible everyday concepts (Burgers et al., 2016). By creating conceptual mappings that emphasize some aspects of issues and hide others (Lakoff and Johnson, 1980), metaphor can affect public attitudes and policy preference (Boeynaems et al., 2017).
@@
We fist identify seven source domains, concepts evoked in discussions of immigration, such as water or vermin. (Figure 1). We use large language models (LLMs) to detect metaphorical words along with document embeddings to detect metaphorical associations in context. Our method requires no manual annotation, but rather just (1) brief concept descriptions, and (2) a handful of example sentences that evoke each metaphorical concept. We evaluate our approach by creating a new crowdsourced dataset of 1.6K tweets labeled for metaphor, and compare several LLMs and prompting strategies. While we focus on U.S. immigration discourse on social media, our approach can be applied to other political, cultural, and discursive contexts.
We then analyze metaphor usage in 400K U.S. tweets about immigration, with a specific focus on the relationship between metaphor, political ideology, and user engagement. We find that conservative ideology is associated with greater use of dehumanizing metaphor, but this effect varies across concepts. Among conservatives, more extreme ideology is associated with higher metaphor use. Surprisingly, while moderate liberals are more likely to use object-related metaphor, extreme liberal ideology is associated with higher use of creature-related metaphor. Moreover, creature-related metaphors are associated with more retweets, and this effect is primarily driven by liberals. We additionally conduct a qualitative analysis to identify diverse contexts in which liberals use such dehumanizing metaphor. Our study reveals nuanced insights only made possible by our novel approach, and highlights the importance and complexity of studying metaphor as a rhetorical strategy in politics. [^1] We will make all annotated data, code, and model outputs available upon publication.
@@ 2 Background
According to Conceptual Metaphor Theory, “the essence of metaphor is understanding and experiencing one thing in terms of another” (Lakoff and Johnson, 1980, p. 5). Metaphor helps people understand complex political issues in terms of more concrete everyday experiences (Burgers et al., 2016). Because metaphors highlight some aspects of issues while downplaying others, they are a type of framing device (Entman, 1993; Burgers et al., 2016) with implications for policy recommendation and political action (Lakoff and Johnson, 1980).
@@
Metaphors of immigrants as animals, vermin, objects, and water have appeared in U.S. immigration debates for centuries (O’Brien, 2003; Card et al., 2022). Metaphorical dehumanization is sometimes overt (e.g. calling immigrants animals) (Santa Ana, 1999), but can be subconscious as some metaphors are conventionalized in immigration discourse (e.g. the water metaphor evoked by waves of immigration) (Porto, 2022). By emphasizing perceived threats from immigrants, dehumanizing metaphors can increase discrimination and harsh immigration policy support (Santa Ana, 1999; Utych, 2018).
@@
Beyond binary ideology, ideology strength may impact metaphor use. From an immigration stance perspective, we would expect the highest use for far-right authors, and lowest for far-left authors. However, both extremes use more negative emotional language than moderates (Alizadeh et al., 2019; Frimer et al., 2019). If metaphors communicate such emotions (Ortony, 1975), their use may be higher for both extremes than for moderates.
@@
How metaphor affects audience’s attitudes and behaviors remains an open question. Critical discourse analysis asserts that discourse is not just shaped by society, but also actively constructs social realities; from this lens, metaphor inherently has strong effects on social and political systems (Charteris-Black, 2006; Boeynaems et al., 2017). However, quantitative experiments have shown mixed (and sometimes irreplicable) results (Thibodeau and Boroditsky, 2011; Steen et al., 2014; Boeynaems et al., 2017; Brugman et al., 2019). Metaphors’ effects are vary across factors such as topic, conceptual domains, message source, political orientation, and personality (Bosman, 1987; Mio, 1997; Robins and Mayer, 2000; Kalmoe, 2014, 2019; Panzeri et al., 2021).
There is experimental evidence of immigration metaphor effects: exposure to animal metaphors increase support for immigration restriction (Utych, 2018), and exposure water increases border wall support (Jimenez et al., 2021). Framing the U.S. as a body amplifies effects of contamination threat exposure on negative attitudes towards immigrants (Zhong and House, 2014). Immigration metaphors’ effects are moderated by contextual variables such as intergroup prejudices and ideology (Marshall and Shapiro, 2018; McCubbins and Ramirez, 2023). For example, animal and vermin metaphors increase support for for-profit immigration detention centers, but only among participants with anti-Latino prejudice (McCubbins and Ramirez, 2023). Due to policy positions and greater sensitivity to threat and disgust (Jost et al., 2003; Inbar et al., 2009), effects may be stronger among conservatives. However, prior work finds that liberals are more susceptible to metaphors’ effects (Thibodeau and Boroditsky, 2011; Hart, 2018).
Moreover, recent work has uncovered resistance to extreme metaphors among conservatives. Republicans are more opposed to immigrant detention centers when exposed to dehumanizing metaphors (McCubbins and Ramirez, 2023). Hart (2021) find that both animal and war metaphors decrease support for anti-immigration sentiments and policies. Boeynaems et al. (2023) show that criminal metaphors of refugees push right-wing populist voters’ opinions away from right-wing immigration stances. Overtly inflammatory metaphors are consciously recognized by audiences, which lead to greater scrutiny of the underlying message (Hart, 2021). Due to liberals’ sensitivity to metaphor and conservatives’ resistance to extreme metaphor, liberals may be more susceptible to the effects of dehumanizing metaphors of immigration.
@@
Our research questions and hypotheses, guided by the aforementioned scholarship, are as follows: [^2] For brevity, we use the term “metaphor” to refer to metaphorical dehumanization of immigrants.
@@
3 Metaphorical concepts
We select seven source domains based on prior literature: animal, vermin, parasite, physical pressure, water, commodity, and war (Table 1). Each concept creates a distinct logic about the perceived threat and plausible remedies. For example, water and physical pressure suggest that immigrants are a threatening force on the host country, metaphorically represented as a container (Charteris-Black, 2006). Potential solutions would reinforce the container, e.g., through border security. The vermin and parasite metaphors make extermination and eradication plausible responses (Steuter and Wills, 2010; Musolff, 2014, 2015).
@@ 4 Measuring Metaphors
We propose a new approach that accounts for both word- and discourse-level metaphor. Even if a message does not borrow specific words from another conceptual domain, its broader logic could still implicitly evoke the metaphor (Brugman et al., 2019). While automatic metaphor processing has primarily focused on lexical information, several computational researchers have argued that broader situational and discourse-level information is necessary to study metaphor production and comprehension (Mu et al., 2019; Dankers et al., 2020). We calculate separate word-level and discourse-level scores using LLMs and document embeddings, respectively, which are then combined to obtain a single metaphor score for each source domain (Figure 2).
@@
We first prompt an LLM to identify and map metaphorical expressions to a source domain (or “none” if no concept applies). For example, in Figure 2, the LLM outputs: {flooding:water}. We test two zero-shot prompts (§A.2). The Simple prompt gives basic instructions and concept names. The Descriptive prompt also provides a definition of metaphor and brief concept descriptions. We evaluate three LLMs: Llama3.1-70B, GPT-4-Turbo-2024-04-09, and GPT-4o-2024-08-06.333All with temperature = 0.
@@
Figure 2: Our method involves calculating the rate of metaphorical words as detected by an LLM (left side), associations between documents and metaphorical concepts with cosine similarity (right side), and adding these measurements together (bottom).
@@
In their study of dehumanizing language, Mendelsohn et al. (2020) estimate associations between target groups and the vermin metaphor as the cosine similarity between word2vec vectors of group labels and an average of vermin-related word embeddings. We adopt this approach with several adjustments. First, we use Transformer-based contextualized embeddings and thus calculate metaphorical associations per document rather than for the entire corpus. Second, we calculate associations between concepts and documents, rather than target group labels. Prior approaches require documents to contain mentions of a target group (Mendelsohn et al., 2020; Card et al., 2022), but many relevant social media posts about sociopolitical issues do not explicitly name the impacted social groups.
We encode each tweet and source domain with SBERT using the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019). Preliminary investigations reveal that just embedding the concept name (e.g. the word water) overemphasizes literal usages (e.g. migrants crossing the sea). To encourage the model to identify metaphorical associations, we represent each concept as a set of “carrier sentences”, which resemble metaphorical usages of the concept but contain little additional semantic information (e.g. they flood in; they hunt them down). We manually construct 104 carrier sentences based on examples from prior literature (Table A3). Each concept has 8-22 carrier sentences (Table A4). We calculate the discourse-level score with respect to a given ${}_{\textsc{concept}}$) as the cosine similarity between the tweet and the average of the carrier sentences’ SBERT representations.
@@ 5 Data
We use the Immigration Tweets Dataset from Mendelsohn et al. (2021), which has 2.6 million English-language tweets from 2018-2019 that contain a keyword related to immigration (e.g. immigration, illegals, undocumented). The dataset does not contain labels for metaphor, but does include inferred labels for several framing typologies. The dataset further provides data about user engagement and authors, including their inferred location at the country level and inferred political ideology based on social network-based models (Compton et al., 2014; Barberá et al., 2015). We select a random sample of 400K tweets whose authors are based in the United States and have an associated ideology estimate and user engagement metrics.
@@
We recruit participants via Prolific to annotate all 1,600 documents. To simplify the task, participants are assigned to one source domain (or the domain-agnostic condition), provided with the relevant codebook portion, and asked to label 20 tweets with respect to their given concept. They are encouraged to make a binary judgment, but can select a third “don’t know” option if needed. [^5] Participants are based in the United States and have completed at least 200 tasks with a $\geq$99approval rate. They are paid \$1.60 per task (\$16/hour). We do not reject any responses through Prolific, but filter out “don’t know” labels and labels from annotators who (1) completed the full task in under three minutes, or (2) gave the same response for all documents.
On average, we obtain eight annotations per tweet, and the mean metaphor score is 0.347 (see §A.1 for more detail about annotators and annotations). The overall inter-annotator agreement is 0.32 (Krippendorff’s $\alpha$) and varies across concepts (Appendix Fig. A1). While lower than between experts, this agreement is both expected and advantageous for our approach. As metaphor comprehension is closely tied to individual cognition, judgments are bound to vary widely across the population. Such heterogeneity reinforces that metaphor is not a clear binary, supporting our continuous measurement approach. Future cognitive science work could investigate why some documents and concepts elicit more disagreement than others.
@@ 6 Model Evaluation
We measure statistical significance with bootstrap tests (n=100). GPT-4o and GPT-4-Turbo with Descriptive prompts and added SBERT outperform all other models, but are not significantly different from each other. We use GPT-4o for large-scale analysis because inference is 4x cheaper.
@@ 7 Analysis
Our analysis uses two sets of regression models. The first set quantifies the role of political ideology and ideological strength in metaphor use (§7.1), and the second set quantifies the association between metaphor, ideology, and user engagement.

### v2

@@ Abstract
Metaphor, discussing one concept in terms of another, is abundant in politics and can shape how people understand important issues. We develop a computational approach to measure metaphorical language, focusing on immigration discourse on social media. Grounded in qualitative social science research, we identify seven source domain concepts evoked in immigration discourse (e.g. water or vermin). We propose and evaluate a novel technique that leverages both word-level and document-level signals to measure metaphor with respect to these source domains. We then study the relationship between metaphor, political ideology, and user engagement in 400K US tweets about immigration. While conservatives tend to use dehumanizing metaphors more than liberals, this effect varies widely across source domains. Moreover, creature-related metaphor is associated with more retweets, especially for liberal authors. Our work highlights the potential for computational methods to complement qualitative approaches in understanding subtle and implicit language in political discourse. [^1] Code and data are available at https://github.com/juliamendelsohn/when_people_are_floods
@@
Julia Mendelsohn University of Maryland juliame@umd.edu                        Ceren Budak University of Michigan cbudak@umich.edu
@@ 1 Introduction
Metaphor, communication of one concept in terms of another, is abundant in political discourse. Rooted in culture and cognition, metaphor structures how we interpret the world (Lakoff and Johnson, 1980), and is deployed consciously and subconsciously to shape our understanding of political issues in terms of accessible everyday concepts (Burgers et al., 2016). By creating conceptual mappings that emphasize some aspects of issues and hide others (Lakoff and Johnson, 1980), metaphor can affect public attitudes and policy preference (Boeynaems et al., 2017).
@@
From prior discourse analysis literature, we first identify seven source domains: concepts evoked in discussions of immigration such as water or vermin (Figure 1). We use large language models (LLMs) to detect metaphorical words along with document embeddings to detect metaphorical associations in context. Our method requires no manual annotation, but rather just (1) brief concept descriptions, and (2) a handful of example sentences that evoke each metaphorical concept. We evaluate our approach by creating a new crowdsourced dataset of 1.6K tweets labeled for metaphor, and compare several LLMs and prompting strategies. While we focus on U.S. immigration discourse on social media, our approach can be applied to other political, cultural, and discursive contexts.
We then analyze metaphor usage in 400K U.S. tweets about immigration, with a specific focus on the relationship between metaphor, political ideology, and user engagement. We find that conservative ideology is associated with greater use of dehumanizing metaphor, but this effect varies across concepts. Among conservatives, more extreme ideology is associated with higher metaphor use. Surprisingly, while moderate liberals are more likely to use object-related metaphor, extreme liberal ideology is associated with higher use of creature-related metaphor. Moreover, creature-related metaphors are associated with more retweets, and this effect is primarily driven by liberals. We additionally conduct a qualitative analysis to identify diverse contexts in which liberals use such dehumanizing metaphor. Our study reveals nuanced insights only made possible by our novel approach, and highlights the importance and complexity of studying metaphor as a rhetorical strategy in politics.
@@ 2 Background
According to Conceptual Metaphor Theory, “the essence of metaphor is understanding and experiencing one thing in terms of another” (Lakoff and Johnson, 1980, p. 5). Metaphor helps people understand complex issues in terms of more concrete everyday experiences (Burgers et al., 2016). Because metaphors highlight some aspects of issues while downplaying others, they are a type of framing device that can affect how people interpret political issues, with implications for policy recommendation and political action (Lakoff and Johnson, 1980; Entman, 1993; Burgers et al., 2016).
@@
There is a large body of critical discourse scholarship focused on the metaphorical framing of immigration, particularly in news media. Metaphors have long been used in communication about immigration as they “facilitate listeners’ grasp of an external, difficult notion of society in terms of a familiar part of life” (Santa Ana, 1999). Metaphors of immigrants as animals, vermin, objects, and water have appeared in U.S. immigration debates for centuries (O’Brien, 2003; Card et al., 2022). Metaphorical dehumanization is sometimes overt (e.g. calling immigrants animals) (Santa Ana, 1999), but can be subconscious as some metaphors are conventionalized in immigration discourse (e.g. the water metaphor evoked by waves of immigration) (Porto, 2022). By emphasizing perceived threats from immigrants, dehumanizing metaphors can increase discrimination and harsh immigration policy support (Santa Ana, 1999; Utych, 2018).
@@
Beyond binary ideology, ideology strength may impact metaphor use. If dehumanizing metaphors are used to communicate threat and advocate for stringent immigration policies, we would expect the highest use for far-right authors. However, both ideological extremes use more negative emotional language than moderates (Alizadeh et al., 2019; Frimer et al., 2019). If metaphors communicate such emotions (Ortony, 1975), their use may be higher for both extremes than for moderates.
@@
How metaphor affects audience’s attitudes and behaviors remains an open question. Critical discourse analysis asserts that discourse is not just shaped by society, but also actively constructs social realities; from this lens, metaphor inherently has strong effects on social and political systems (Charteris-Black, 2006; Boeynaems et al., 2017). However, quantitative experiments have shown mixed (and sometimes irreplicable) results (Thibodeau and Boroditsky, 2011; Steen et al., 2014; Boeynaems et al., 2017; Brugman et al., 2019). Metaphors’ effects vary across factors such as topic, conceptual domains, message source, political orientation, and personality (Bosman, 1987; Mio, 1997; Robins and Mayer, 2000; Kalmoe, 2014, 2019; Panzeri et al., 2021).
There is experimental evidence of immigration metaphor effects: exposure to animal metaphors increases support for immigration restriction, with the effect mediated by emotions of anger and disgust (Utych, 2018), and exposure to the water metaphor increases border wall support (Jimenez et al., 2021). Framing the U.S. as a body amplifies effects of contamination threat exposure on negative attitudes towards immigrants (Zhong and House, 2014). Immigration metaphors’ effects are moderated by contextual variables such as intergroup prejudices and ideology (Marshall and Shapiro, 2018; McCubbins and Ramirez, 2023). For example, animal and vermin metaphors increase support for for-profit immigration detention centers, but only among participants with anti-Latino prejudice (McCubbins and Ramirez, 2023). Political ideology may also moderate metaphorical framing effects. Due to policy positions and greater sensitivity to threat and disgust (Jost et al., 2003; Inbar et al., 2009), effects may be stronger among conservatives. However, prior work finds that liberals are more susceptible to metaphors’ effects (Thibodeau and Boroditsky, 2011; Hart, 2018). Moreover, Lahav and Courtemanche (2012) find that threat framing has stronger effects on immigration policy attitudes for liberals compared to conservatives. Hart (2018) makes an argument for entrenchment: conservatives have more fixed attitudes and opinions, “while liberals formulate their views on a more context-dependent basis taking into account local information supplied by texts.”
Additionally, recent work has uncovered resistance to extreme metaphors among conservatives. Republicans are more opposed to immigrant detention centers when exposed to dehumanizing metaphors (McCubbins and Ramirez, 2023). Hart (2021) find that both animal and war metaphors decrease support for anti-immigration sentiments and policies. Boeynaems et al. (2023) show that criminal metaphors of refugees push right-wing populist voters’ opinions away from right-wing immigration stances. Overtly inflammatory metaphors are consciously recognized by audiences, which lead to greater scrutiny of the underlying message (Hart, 2021). Due to liberals’ sensitivity to metaphor and conservatives’ resistance to extreme metaphor, liberals may be more susceptible to the effects of dehumanizing metaphors of immigration. Our work investigates if metaphorical framing effects occur outside of lab experiments by considering user engagement with real-world messages on social media.
@@
Guided by the aforementioned scholarship, we put forth four research questions and hypotheses:  [^2] For brevity, we use the term “metaphor” to refer to metaphorical dehumanization of immigrants.
@@
3 Source domains
We select seven source domains that have been well-documented in prior discourse analysis literature about immigration: animal, vermin, parasite, physical pressure, water, commodity, and war (Table 1). While they all liken immigrants to non-human entities, each source domain creates a distinct logic about the perceived threat and plausible remedies. For example, water and physical pressure suggest that immigrants are a threatening force on the host country, metaphorically represented as a container (Charteris-Black, 2006). Potential solutions would reinforce the container, e.g., through border security. Organism-related metaphors such as vermin and parasite create conceptual mappings through which the existence of immigrants is a threat, thus making extermination and eradication plausible responses (Steuter and Wills, 2010; Musolff, 2014, 2015).
Some source domains may be seen as more extreme and blatantly dehumanizing than others (e.g. vermin vs. commodity). There is also heterogeneity within each source domain; for example, historical waves of immigration and immigrants are flooding in both evoke the water metaphor, but the former may be viewed as more conventional with a more neutral valence. Note that there are relevant source domains for immigration discourse beyond these seven, such as plants (Santa Ana, 1999; Gonçalves, 2024), which could be further explored in future work.
@@ 4 Measuring Metaphors
We propose a new approach that accounts for both word- and discourse-level metaphor. Even if a message does not borrow specific words from another source domain, its broader logic could still implicitly be evoked (Brugman et al., 2019). While automatic metaphor processing has primarily focused on lexical information, several computational researchers have argued that broader situational and discourse-level information is necessary to study metaphor production and comprehension (Mu et al., 2019; Dankers et al., 2020). We calculate separate word-level and discourse-level measurements using LLMs and document embeddings, respectively, which are then combined to obtain a single metaphor score for each source domain (Figure 2).
@@
We first prompt an LLM to identify and map metaphorical expressions to a source domain (or “none” if no source domain applies). For example, in Figure 2, the LLM outputs: {flooding:water}. We test two zero-shot prompts. The Simple prompt gives basic instructions and concept names. The Descriptive prompt also provides a definition of metaphor and brief concept descriptions. We evaluate three LLMs: Llama3.1-70B, GPT-4-Turbo-2024-04-09, and GPT-4o-2024-08-06.333All with temperature = 0. See Appendix for full prompts.
@@
Figure 2: Our method involves calculating the rate of metaphorical words as detected by an LLM (left side), metaphorical associations between documents and source domain concepts with cosine similarity (right side), and adding these measurements together (bottom).
@@
In their study of dehumanizing language, Mendelsohn et al. (2020) estimate associations between target groups and the vermin metaphor as the cosine similarity between word2vec representations of group labels and an average of vermin-related word embeddings. We adopt this approach with several adjustments. First, we use Transformer-based contextualized embeddings instead of static word embeddings. Second, we calculate associations between source domains and documents, rather than target group labels. Prior approaches require documents to contain mentions of a target group (Mendelsohn et al., 2020; Card et al., 2022), but many relevant social media posts about sociopolitical issues do not explicitly name the impacted social groups.
[…]
@@ 5 Data
We use the Immigration Tweets Dataset from Mendelsohn et al. (2021), which has 2.6 million English-language tweets from 2018-2019 that contain a keyword related to immigration (e.g. immigration, illegals, undocumented). The dataset does not contain labels for metaphor, but does include inferred labels for issue-generic, issue-specific, and episodic and thematic frames. The dataset further provides data about user engagement and authors, including their inferred location at the country level and inferred political ideology based on social network-based models (Compton et al., 2014; Barberá et al., 2015). We select a random sample of 400K tweets whose authors are based in the United States and have an associated ideology estimate and user engagement metrics.
The Immigration Tweets dataset already includes estimates of authors’ political ideology: Mendelsohn et al. (2021) use the Bayesian Spatial Following model from Barberá et al. (2015). Leveraging assumptions of homophily, this model estimates ideology on a continuous scale based on the Twitter following network and does not consider text. The resulting ideal points are continuous estimates ranging from approximately -3 (most liberal) to +3 (most conservative), with values near 0 representing political moderates. We construct two variables to separately examine the effects of liberal vs. conservative ideology (sign of the ideal point) and ideological strength (magnitude of the ideal point).
@@
We recruit participants via Prolific to annotate all 1,600 documents. To simplify the task, participants are assigned to one source domain (or the domain-agnostic condition), provided with the relevant codebook portion, and asked to label 20 tweets with respect to their given source domain. They are encouraged to make a binary judgment, but can select a third “don’t know” option if needed. [^5] Participants are based in the United States and have completed at least 200 tasks with a $\geq$99approval rate. They are paid \$1.60 per task (\$16/hour). We do not reject any responses through Prolific, but filter out “don’t know” labels and labels from annotators who (1) completed the full task in under three minutes, or (2) gave the same response for all documents.
On average, we obtain eight annotations per tweet, and the mean metaphor score is 0.347 (see §A.1 for details about annotators and annotations). The overall inter-annotator agreement is 0.32 (Krippendorff’s $\alpha$) and varies across source domains (Fig. A1). While lower than between experts, this agreement is both expected and advantageous for our approach. As metaphor comprehension is closely tied to culture and individual cognition, judgments may vary widely across the population. Such heterogeneity reinforces that metaphor is not a clear binary, supporting our continuous measurement approach. Future cognitive science work could investigate why some documents and source domains elicit more disagreement than others.
@@ 6 Model Evaluation
We measure statistical significance with bootstrap tests (n=100) with a 95% confidence interval. GPT-4o and GPT-4-Turbo with Descriptive prompts and added SBERT outperform all other models, but are not significantly different from each other. We use GPT-4o for large-scale analysis because inference is 4x cheaper.
@@ 7 Analysis
Our analysis uses two sets of regression models. The first quantifies the role of political ideology and ideological strength in metaphor use (§7.1), and the second quantifies the association between metaphor, ideology, and user engagement (§7.2).